ACTIVITY:
The average cholesterol content of a certain canned good is 215 mg, and the standard
deviation is 15 mg. Assume that the variable is normally distributed. If a sample of 25 canned goods
are selected, what is the probability that the mean of the sample will be larger than 220 mg
1
Expert's answer
2021-06-01T09:10:00-0400

Let "X=" the average cholesterol content of a certain can goods in milligrams.


"X\sim N(\mu, \sigma^2/n)\\\"

Then "Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0. 1)"


Given that,

"X= 220\\\mu=215\\\sigma = 15\\n=25"


"z=\dfrac{X-\mu}{\sigma/\sqrt n}=\dfrac{220-215}{15/\sqrt{25}}=\dfrac{25}{15}=1.67"


Thus, probability that mean of sample is greater than 220 mg is

"P(z>220)=P(z>1.67)"


So, from standard normal distribution table:





"P(z>1.67)=0.0475"


Hence, The probability that mean of sample is greater than 220 mg is 0.0475